OpenStudy (brittany16):

Solve the equation. -3f - 14 = 1

OpenStudy (nubeer):

-5

OpenStudy (kamille):

-3f - 14 = 1 -3f=14+1 -3f=15|:(-3) f=-5

OpenStudy (anonymous):

-3f = 1 +14 -3f = 15 f = 15 / -3 f = -5

OpenStudy (anonymous):

http://i.imgur.com/UUfF1.png